Compressor monitor system (CMS)

Links :
06jun11: testing the compressor monitor system on the telescope.
14apr11: testing the compressor monitor system in the lab

06jun11: on telescope testing of CMS. (top)

    The 327 vlbi run on 05jun11 reported a 1 Mhz comb  in the data. On 06jun11 we did some tests to see if the  problem was coming from the compressor monitor system (which had been turned on during the 05jun11).
The setup was:
Processing the data:

The image shows the dynamic spectra for the 750 seconds of data (.gif):
The plot shows  average spectrum with the compressor monitor on and off (.ps) (.pdf):


processing: x101/110606/

14/15apr11:Compressor monitoring system in the lab (top)

    The Compressor monitoring system was tested for rfi by luis Quintero 14apr11 in the screened room. It was retested 15apr11 after some added shielding.The setup was:
The plots show rfi from the compressor monitor system:


processing: x101/110415/